A 50m by 20m rectangular lawn is mowed by cuting a uniform strip along it's perimeter. This leaves 4/5 of the lawn area uncut. Determing the width of the cutto the nearest hundredth of a meter
Answer by scott8148(6628)

(50 - 2w)(20 - 2w) = 800 ___ (25 - w)(10 - w) = 400

w^2 - 35w - 150 = 0

use the quadratic formula to find w